# One-Embedding-Fits-All: Efficient Zero-Shot Time Series Forecasting by a Model Zoo

### v1

@@ 4.1 Advantage Subset Characterization
4.2 Model-task Representations Co-Embedding
Even with model-specific subsets, another major challenge is embedding tasks and models into a common space despite the non-stationary and multi-channel nature of time series. We address this with a co-embedding extractor that maps both into a unified semantic space for efficient and consistent similarity-based matching.
@@ 4.2.1 Build Co-Embedding Extractor
To facilitate efficient model selection, we introduce a general co-embedding extractor $\psi:\mathbb{R}^{T}\rightarrow\mathbb{R}^{D}$, which maps both target time series and foundation models into a shared vector space. The extractor is trained on an independent dataset $\mathcal{D}^{*}$ (disjoint from both zoo pretraining corpora and downstream forecasting tasks) to prevent information leakage, with an encoder producing latent representations and a decoder reconstructing the original series. The complete training objective integrates three components (detailed formulations are provided in Appendix C.1):
@@
Here, $\mathcal{L}_{\text{Reconstruction}}$ encourages temporal fidelity, $\mathcal{L}_{\text{Constraint}}$ applies masked-view contrastive learning to enhance robustness, and $\mathcal{L}_{\text{Transfer}}$ introduces a novel transferability alignment specifically designed for time-series tasks, supervising cross-task similarity with $1$-MSE scores. The coefficient $\lambda$ is a balancing factor that controls the relative contribution of the transferability loss to the overall objective. This transferability loss is the core innovation that ensures the learned space reflects actual generalization ability across datasets. Once trained, the extractor provides the foundation for selecting the most suitable models for new tasks, facilitating zero-shot forecasting with high efficiency. The detailed formulations, training setup, and dataset specifications are provided in Appendix C.1.
@@ 4.2.2 Model-task Representation Construction
For model representation, each foundation model $\phi_{m}$ is encoded through its advantage subset $\mathcal{D}_{m}^{\text{adv}}$. The extractor $\psi$ encodes all samples $\mathbf{x}_{i}\in\mathcal{D}_{m}^{\text{adv}}$ into embeddings, which are then averaged into a single representation vector $\mathbf{r}_{m}\in\mathbb{R}^{D}$. Collecting these vectors across all $M$ models yields the model zoo representation library:
$\mathbf{R}_{\text{zoo}}=\left[\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{M}\right]^{\top}\in\mathbb{R}^{M\times D}.$ (5)
This shared library enables efficient similarity computation and supports seamless zoo expansion: new models can be integrated by appending their representations without recomputing existing entries(detailed in the Appendix D.2), preserving the one-embedding-fits-all principle of ZooCast.
For the target series $X=\{\mathbf{x}_{c}\}_{c=1}^{C}$ with $C$ channels, task representations are constructed by sampling segments of length $T$ from each channel. The extractor $\psi$ maps these segments into embeddings:
$\mu_{c}\;=\;\psi(\mathbf{x}_{c})\in\mathbb{R}^{D},\quad c=1,\dots,C,\qquad\mu\;=\;[\mu_{1},\mu_{2},\dots,\mu_{C}]^{\top}\in\mathbb{R}^{C\times D}.$ (6)
This co-embedding mechanism establishes a unified vector space where both models and tasks reside, effectively transforming model selection into a lightweight similarity measurement between $\mathbf{R}_{\text{zoo}}$ and $\mu$. The shared representation space preserves semantic relationships between model capabilities and task requirements while enabling $\mathcal{O}(1)$ time complexity for matching operations.
@@ 4.3 Matching TSFMs for Target Time-Series
Model-task Similarity Computation. Given a target series $X\in\mathbb{R}^{C\times T}$ with $C$ channels, we compute channel-wise similarity scores between $\mu_{c}$ and all model representations $\mathbf{r}_{m}\in\mathbf{R}_{\text{zoo}}$. Each similarity score is modulated by a model-specific weight $w_{m}$ (derived from the advantage subset; see Appendix D.1 for details):
$\text{sim}_{m,c}\;=\;w_{m}\cdot\frac{\mathbf{r}_{m}^{\top}\mu_{c}}{\|\mathbf{r}_{m}\|\,\|\mu_{c}\|},\quad\mathbf{A}=[\text{sim}_{m,c}]_{C\times M}.$ (7)
Since both model and task representations are low-dimensional vectors ($D\leq 128$), this computation remains efficient even for large model zoos ($M>100$) and high-channel series ($C>50$), typically completing within one second.
Error-Correcting Consensus Ranking. To aggregate channel-specific recommendations into a unified ranking, we employ an error-correcting code framework. First, encode each channel’s top-$r$ models ($r=3$) into a binary matrix $\mathbf{B}\in\{0,1\}^{C\times M}$, where $B_{c,m}=1$ indicates model $\phi_{m}$ is recommended for channel $c$. Then compute Hamming distances to the ideal consensus:
$h_{m}\;=\;\sum_{c=1}^{C}\mathbb{I}(B_{c,m}=0),\quad\mathbf{r}_{\text{final}}\;=\;\mathop{\mathrm{argsort}}_{m}h_{m}.$ (8)
This ECOC ECOC -inspired approach mimics error correction in digital communications, where cross-channel redundancies mitigate individual channel noise. The final ranking $\mathbf{r}_{\text{final}}$ prioritizes models with broad consensus across channels.
@@ 5 Experiments
We comprehensively evaluate ZooCast through three key analyses: (1) zero-shot point forecasting performance against baseline TSFMs on Gift-Eval GE , a large-scale and best-recognized zero-shot benchmark. (Section 5.1), (2) test-time scaling behavior of ensembles guided by ZooCast’s model rankings and ablation studies on critical parameters when constructing the advantage subset in ZooCast (Section 5.2), (3) computational efficiency analysis of repr-based model zoo zero-shot forecasting comparison with forward all TSFMs (Section 5.3). All experiments were conducted under identical hardware and evaluation protocols to ensure fair comparison.
@@ 5.1 Evaluation on GIFT-Eval benchmark
Table 3: As most existing model selection methods do not support time-series or regression tasks, and LogME LogME was originally designed for time-series forecasting, we adapt its workflow and restrict its application to univariate forecasting, constructing LogME-guided (L.M.) ensembles. The evaluation datasets are reduced to 32 univariate configurations, with all other settings identical to Table 2.
@@
Setup. Following previous work GE ; liu2025sundialfamilyhighlycapable , we adopt sMAPE as the evaluation metric, excluding CRPS since ZooCast targets point forecasting, whereas CRPS targets probabilistic settings and would bias evaluation of deterministic outputs. Compared to MAPE, sMAPE is symmetric and more robust to small or zero values. We report raw sMAPE values without normalization by a Seasonal Naive baseline. To ensure fair comparison across diverse datasets, we compute a Rank score by ranking models within each configuration based on sMAPE and then averaging ranks across all datasets, which mitigates scale differences. Ensemble results are obtained via the naive strategy of simply averaging predictions of the selected TSFMs, highlighting the effectiveness of our framework.
Baseline. We evaluate 13 TSFMs as single model prediction baselines: Chronos (Chr.bT/bM/bS/bB) chronos , Moirai (Moi.S/B/L) MOIRAI , VisionTS (Vis.B/L/H) VisionTS , TimesFM (TFM.1/2) TimesFM , and Sundial (Sun.B) liu2025sundialfamilyhighlycapable . All models are tested in their original zero-shot configurations. In Table 2 and 3, we report two ensemble baselines: the naive ensemble of all TSFMs (All-13) and LogME-guided ensembles (L.M.), which also serve as our model selection baseline. In Figure 4, we additionally include four classical model selection strategies as baselines: Recent Selection, All Current Ensemble, Latest Selection, and Current Best Selection. More experimental settings are provided in Appendix C.2.
@@
Results. Table 2 and 3 evaluates the most comprehensive model zoo at a single time point, reporting aggregated performance across 97 configurations from 23 datasets. ZooCast-guided ensembles consistently outperform all baselines, with Top-5 Z.C.En. and Top-3 Z.C.En. both delivering highly competitive Rank scores, demonstrating stable average superiority across diverse datasets. Since sMAPE can be affected by extreme values on certain datasets, the clear advantage in Rank provides stronger evidence of overall robustness. Figure 4 further reports performance dynamics under sequential model releases: as more recent TSFMs are added, the sMAPE of Top-3 Z.C.En. steadily improves over time, while its Rank remains consistently the best. This indicates that the larger and more up-to-date the model zoo becomes, the more stable ZooCast’s selection advantage is.
@@
5.2 Ensemble Scaling Behavior of ZooCast
The test-time scaling law scaling_law refers to the principle that by increasing the prediction stages or computational cost, the prediction performance can improve. A robust scaling strategy should enhance performance with minimal additional computational expense compared to single-model predictions. In ensemble strategies, the optimal approach is to achieve this improvement by accurately selecting a small number of models, avoiding excessive computational costs.
ZooCast provides an effective solution by leveraging model ranks to guide ensemble selection. By choosing only a few top-ranked models, it achieves strong predictive performance at a fraction of the cost of evaluating all candidates. This effect is clearly illustrated in Figure 5, which demonstrates the ensemble scaling behavior as performance steadily improves with a small number of selected models. Furthermore, Figures 5(a) and 5(b) validate parameter sensitivity by varying the initial settings (size=1000, threshold=1) from Section 5.1. Across all tested configurations, the Top-3 ensemble consistently outperforms both the naive ensemble (All-13 En.) and the strongest single model (Chr.bB), confirming the robustness of ZooCast under different parameter choices.
@@ 5.3 Efficiency Analysis
Results. Table 4 presents the runtime comparison across three execution phases. The Full Forward phase simulates a brute-force baseline where all TSFMs are evaluated on every new forecasting task, resulting in a total cost of 88024 seconds. This represents the standard zero-shot approach with no reuse of prior computation. In contrast, the Precompute phase incurs a one-time cost of 123 seconds for building the model repr. via forward passes on the zoo repr. set $\mathcal{D}$, excluded from speedup calculations. The Selection phase uses the pre-built zoo repr. library $\mathbf{R}_{\text{zoo}}$ for lightweight model selection, taking 1024 seconds across all tasks. The Forecast phase directly generates predictions based on the selection results. Interestingly, we find that the actual speedup achieved by ZooCast is far greater than the number of models in the zoo. This is because the models selected by ZooCast are not only more accurate but also faster to run. While this may partly be coincidental, it suggests that ZooCast tends to favor models that are both superior in performance and more efficient, thereby achieving a dramatic speedup over full forward evaluation.
Specifically, a single matrix computation between $\mathbf{R}_{\text{zoo}}$ and the task representation $\mu_{c}$ yields the recommendation. This approach enables a constant-time $\mathcal{O}(1)$ process, bypassing the need for any forward computations. By relying on lightweight precomputed representations, we reduce selection complexity from $\mathcal{O}(MN)$ to $\mathcal{O}(1)$ per task, allowing for instant recommendations at inference time and making this repr-based selection highly practical for large-scale zero-shot forecasting. Detailed computational complexity and memory usage analysis is provided in the Appendix D.3.
@@ Appendix A Mathematical Notations
$X\in\mathbb{R}^{C\times T}$ Multivariate input time series with $C$ channels and length $T$
$\mathbf{x}_{t}\in\mathbb{R}^{C}$ Observation vector at time step $t$ across all channels
$\mathbf{x}_{c}\in\mathbb{R}^{T}$ The $c$-th channel sequence of length $T$
$Y\in\mathbb{R}^{C\times H}$ Ground-truth future multivariate sequence of horizon $H$
$\hat{Y}\in\mathbb{R}^{C\times H}$ Predicted future multivariate sequence of horizon $H$
$\phi_{m}$ The $m$-th foundation model (TSFM) in the zoo, $m=1,\dots,M$
@@
$\mathcal{D}$ Zoo characterization set containing $n$ sampled subsequences
$\mathbf{x}_{i}$ A subsequence sampled from $\mathcal{D}$
$\mathbf{E}\in\mathbb{R}^{M\times n}$ Precomputed MSE/error matrix on $\mathcal{D}$
$\mathcal{D}_{m}^{\text{adv}}$ Advantage subset of model $\phi_{m}$, defined by its relative performance
$d_{m}$ Cardinality of $\mathcal{D}_{m}^{\text{adv}}$ above threshold
$w_{m}$ Weight of model $\phi_{m}$, derived from $d_{m}$
$\sigma_{i}$ Inter-model variance of errors on sample $\mathbf{x}_{i}$
@@
$\mathbf{A}\in\mathbb{R}^{C\times M}$ Similarity matrix between channels and models
$\mu_{c}\in\mathbb{R}^{D}$ Embedding of channel $c$, $\mu_{c}=\psi(\mathbf{x}_{c})$
$\mathbf{R}_{\text{zoo}}\in\mathbb{R}^{M\times D}$ Model zoo representation library of all TSFM embeddings
@@
$\mathbf{r}_{\text{final}}$ Final model ranking after error-correcting consensus
$\Delta P$ Performance gain relative to the best single model
$\eta$ Selection efficiency, defined as accuracy–cost tradeoff ratio
$T_{\text{select}},T_{\text{infer}}$ Time costs of model selection and inference, respectively

### v2

@@ 4.2.1 Build Co-Embedding Extractor
To facilitate efficient model selection, we introduce a general co-embedding extractor $\psi:\mathbb{R}^{T}\to\mathbb{R}^{D}$, which maps both target time series and foundation models into a shared vector space. The extractor is trained on an independent dataset $\mathcal{D}^{*}$ (disjoint from both zoo pretraining corpora and downstream forecasting tasks) to prevent information leakage, with an encoder producing latent representations and a decoder reconstructing the original series. The complete training objective integrates three components (detailed formulations are provided in Appendix E.2):
@@
Here, $\mathcal{L}_{\text{Reconstruction}}$ encourages temporal fidelity, $\mathcal{L}_{\text{Constraint}}$ applies masked-view contrastive learning to enhance robustness, and $\mathcal{L}_{\text{Transfer}}$ introduces a novel transferability alignment specifically designed for time-series tasks, supervising cross-task similarity with $1$-MSE scores. The coefficient $\lambda$ is a balancing factor that controls the relative contribution of the transferability loss to the overall objective. This transferability loss is the core innovation that ensures the learned space reflects actual generalization ability across datasets. Once trained, the extractor provides the foundation for selecting the most suitable models for new tasks, facilitating zero-shot forecasting with high efficiency. The detailed formulations, training setup, and dataset specifications are provided in Appendix E.2.
@@ 4.2.2 Model-task Representation Construction
For model representation, each foundation model $\phi_{m}$ is encoded through its advantage subset $\mathcal{D}_{m}^{\text{adv}}$, where all subsequences are embedded and then averaged into a single vector ${\bm{r}}_{m}$. For the target series ${\bm{X}}\in\mathbb{R}^{C\times T}$, task representations are constructed by sampling segments of length $t$ from each channel. The extractor $\psi$ maps these segments into embeddings:
${\bm{r}}_{m}\;=\;\tfrac{1}{|\mathcal{D}_{m}^{\text{adv}}|}\sum_{{\bm{x}}_{i}\in\mathcal{D}_{m}^{\text{adv}}}\psi({\bm{x}}_{i})\in\mathbb{R}^{D},\quad m=1,\dots,M,\qquad{\bm{R}}_{\text{zoo}}=[{\bm{r}}_{1},\dots,{\bm{r}}_{M}]^{\top}\in\mathbb{R}^{M\times D}.$
${\bm{\mu}}_{c}\;=\;\psi({\bm{x}}_{c})\in\mathbb{R}^{D},\quad c=1,\dots,C,\qquad{\bm{R}}_{\text{task}}=[{\bm{\mu}}_{1},\dots,{\bm{\mu}}_{C}]^{\top}\in\mathbb{R}^{C\times D}.$
Collecting ${\bm{r}}_{m}$ across all $M$ models yields the model zoo representation library ${\bm{R}}_{\text{zoo}}$, while ${\bm{R}}_{\text{task}}$ serves as the task embedding matrix. This symmetric formulation highlights that model representations are averaged over advantage subsets, whereas task representations are constructed channel-wise from sampled segments. Together, they enable efficient similarity computation in a unified vector space, supporting seamless zoo expansion (Appendix F.2) and a lightweight similarity measurement between ${\bm{R}}_{\text{zoo}}$ and ${\bm{R}}_{\text{task}}$, with $\mathcal{O}(1)$ time complexity for matching operations.
@@ 4.3 Matching TSFMs for Target Time-Series
For a target series $\bm{X}$}, we compute the channel–model similarity score $\text{sim}_{m,c}$ between each channel embedding ${\bm{\mu}}_{c}\in\mathbb{R}^{D}$ and each model representation ${\bm{r}}_{m}\in{\bm{R}}_{\text{zoo}}$, weighted by a model-specific reliability factor $w_{m}$ (see Appendix F.1 for details). For each channel, the top-$r$ models ($r=3$) with the largest $\text{sim}_{m,c}$ are marked in a binary recommendation matrix ${\bm{B}}\in\{0,1\}^{C\times M}$, where $B_{c,m}=1$ indicates model $\phi_{m}$ is recommended for channel $c$. Aggregating these votes through Hamming distance $h_{m}$ yields the final ranking ${\bm{r}}_{\text{final}}$:
$\text{sim}_{m,c}=w_{m}\cdot\frac{{\bm{r}}_{m}^{\top}{\bm{\mu}}_{c}}{\|{\bm{r}}_{m}\|\|{\bm{\mu}}_{c}\|},\quad h_{m}=\sum_{c=1}^{C}\mathbb{I}(B_{c,m}=0),\quad{\bm{r}}_{\text{final}}=\mathop{\mathrm{argsort}}_{m}h_{m}.$ (5)
This ECOC (Kong & Dietterich, 1995)-inspired approach mimics error correction in digital communications, where cross-channel redundancies mitigate individual channel noise. The final ranking ${\bm{r}}_{\text{final}}$ prioritizes models with broad consensus across channels.
@@ 5 Experiments
We comprehensively evaluate ZooCast through three key analyses: (1) zero-shot point forecasting performance against baseline TSFMs on Gift-Eval (Aksu et al., 2024), a large-scale and best-recognized zero-shot benchmark. (Section 5.1), (2) test-time scaling behavior of ensembles guided by ZooCast’s model rankings and ablation studies on critical parameters when constructing the advantage subset in ZooCast (Section 5.2), (3) computational efficiency analysis of repr-based model zoo zero-shot forecasting comparison with forward all TSFMs (Section 5.3). All experiments were conducted under identical hardware and evaluation protocols to ensure fair comparison.
@@ 5.1 Evaluation on GIFT-Eval benchmark
Table 3: As most existing model selection methods do not support time-series or regression tasks, and LogME was originally designed for time-series forecasting, we adapt its workflow and restrict its application to univariate forecasting, constructing LogME-guided (L.M.) ensembles. The evaluation datasets are reduced to 32 univariate configurations, with all other settings identical to Table 2.
@@
Setup. Following previous work (Aksu et al., 2024; Liu et al., 2025), we adopt sMAPE as the evaluation metric, excluding CRPS since ZooCast targets point forecasting, whereas CRPS targets probabilistic settings and would bias evaluation of deterministic outputs. Compared to MAPE, sMAPE is symmetric and more robust to small or zero values. We report raw sMAPE values without normalization by a Seasonal Naive baseline. To ensure fair comparison across diverse datasets, we compute a Rank score by ranking models within each configuration based on sMAPE and then averaging ranks across all datasets, which mitigates scale differences. Ensemble results are obtained via the naive strategy of simply averaging predictions of the selected TSFMs, highlighting the effectiveness of our framework.
Baseline. We evaluate 13 TSFMs as single model prediction baselines: Chronos (Chr.bT/bM/bS/bB) (Ansari et al., 2024), Moirai (Moi.S/B/L) (Woo et al., 2024), VisionTS (Vis.B/L/H) (Chen et al., 2025), TimesFM (TFM.1/2) (Das et al., 2024), and Sundial (Sun.B) (Liu et al., 2025). All models are tested in their original zero-shot configurations. In Table 2 and 3, we report two ensemble baselines: the naive ensemble of all TSFMs (All-13) and LogME-guided ensembles (L.M.), which also serve as our model selection baseline. In Figure 4, we additionally include four classical model selection strategies as baselines: Random Selection, All Current Ensemble, Latest Selection, and Current Best Selection. More experimental settings are provided in Appendix D.
@@
Results. Table 2 and 3 report aggregated performance across 97 configurations from 23 datasets. ZooCast-guided ensembles consistently outperform all baselines, with Top-5 Z.C.En. and Top-3 Z.C.En. both delivering highly competitive Rank scores, demonstrating stable average superiority across diverse datasets. Since sMAPE can be affected by extreme values on certain datasets, the clear advantage in Rank provides stronger evidence of overall robustness. Figure 4 further reports performance dynamics under sequential model releases: as more recent TSFMs are added, the sMAPE of Top-3 Z.C.En. steadily improves over time, while its Rank remains consistently the best. This indicates that the larger and more up-to-date the model zoo becomes, the more stable ZooCast’s selection advantage is.
@@
5.2 Ensemble Scaling Behavior of ZooCast
The test-time scaling law (Zhang et al., 2025a) refers to the principle that by increasing the prediction stages or computational cost, the prediction performance can improve. A robust scaling strategy should enhance performance with minimal additional computational expense compared to single-model predictions. In ensemble strategies, the optimal approach is to achieve this improvement by accurately selecting a small number of models, avoiding excessive computational costs.
ZooCast provides an effective solution by leveraging model ranks to guide ensemble selection. By choosing only a few top-ranked models, it achieves strong predictive performance at a fraction of the cost of evaluating all candidates. This effect is clearly illustrated in Figure 5, which demonstrates the ensemble scaling behavior as performance steadily improves with a small number of selected models. Furthermore, Figures 5(a) and 5(b) validate parameter sensitivity by varying the initial settings (size=1000, threshold=1) from Section 5.1. Across all tested configurations, the Top-3 ensemble consistently outperforms both the naive ensemble (All-13 En.) and the strongest single model (Sun.B), confirming the robustness of ZooCast under different parameter choices.
@@ 5.3 Efficiency Analysis
Results. Table 4 presents the runtime comparison across three execution phases. The Full Forward phase simulates a brute-force baseline where all TSFMs are evaluated on every new forecasting task, resulting in a total cost of 88024 seconds. This represents the standard zero-shot approach with no reuse of prior computation. In contrast, the Precompute phase incurs a one-time cost of 123 seconds in total for building the representation of each model via forward passes on the zoo representation set $\mathcal{D}$. The Selection phase uses the pre-built zoo representation library ${\bm{R}}_{\text{zoo}}$ for lightweight model selection, taking 1024 seconds across all tasks. The Forecast phase directly generates predictions based on the selection results. Interestingly, we find that the actual speedup achieved by ZooCast is far greater than the number of models in the zoo. This is because the models selected by ZooCast are not only more accurate but also faster to run. While this may partly be coincidental, it suggests that ZooCast tends to favor models that are both superior in performance and more efficient, thereby achieving a dramatic speedup over full forward evaluation.
Specifically, a single matrix computation between ${\bm{R}}_{\text{zoo}}$ and the task representation ${\bm{R}}_{\text{task}}$ yields the recommendation. This approach enables a constant-time $\mathcal{O}(1)$ process, bypassing the need for any forward computations. By relying on lightweight precomputed representations, we reduce selection complexity from $\mathcal{O}(MN)$ to $\mathcal{O}(1)$ per task, allowing for instant recommendations at inference time and making this repr-based selection highly practical for large-scale zero-shot forecasting. Detailed computational complexity and memory usage analysis is provided in the Appendix F.3.
@@ 6 Conclusion
Reproducibility Statement
The results presented in this paper are fully reproducible. Our work focuses on proposing a novel model selection framework for existing time series foundation models (TSFMs), all of which have publicly available implementations and evaluation pipelines documented in the GIFT-Eval benchmark. The minor adaptations we introduce for reproducibility are described in detail in Appendix D. The datasets used in all experiments, along with their preprocessing protocols, are entirely drawn from the publicly available GIFT-Eval benchmark, ensuring transparent and standardized evaluation. For the efficiency analysis, a detailed explanation of the theoretical complexity is provided in Appendix F.3. Moreover, the design of the proposed selection framework, including the advantage subset construction, co-embedding extractor, and consensus ranking mechanism, is fully described in both the main text and the appendix to ensure clarity. To further support reproducibility, the complete source code will be released as supplemental material upon the paper’s acceptance.
@@ Appendix A Mathematical Notations
We summarize the mathematical notations used throughout the paper in Table 5. This convention ensures consistency and aligns with the standardized notation file in the ICLR template.
@@
${\bm{X}}\in\mathbb{R}^{C\times T}$ Multivariate input time series with $C$ channels and length $T$
${\bm{x}}_{t}\in\mathbb{R}^{C}$, ${\bm{x}}_{c}\in\mathbb{R}^{T}$ Observation at time step $t$; the $c$-th channel sequence of length $T$
${\bm{Y}},\hat{{\bm{Y}}}\in\mathbb{R}^{C\times H}$ Ground-truth and predicted multivariate sequences of horizon $H$
$\mathcal{Z}=\{\phi_{1},\dots,\phi_{M}\}$ Model zoo of $M$ foundation models (TSFMs)
${\bm{r}}_{\text{final}}$ Final model ranking after error-correcting consensus
$\Delta P$ Performance gain relative to the best single model
$\eta$ Selection efficiency, defined as accuracy–cost tradeoff ratio
$T_{\text{all}}$ Total runtime including both model selection and inference
@@
$\mathcal{D}$, ${\bm{x}}_{i}$ Zoo characterization set containing $n$ sampled subsequence ${\bm{x}}_{i}$
${\bm{E}}\in\mathbb{R}^{M\times n}$ Precomputed MSE/error matrix on $\mathcal{D}$
$s_{m,i},\tau$ Advantage score of model $\phi_{m}$ on sample ${\bm{x}}_{i}$ and its threshold for subset construction
$\mathcal{D}^{\mathrm{adv}}_{m}$, $d_{m}$, $w_{m}$ Advantage subset of $\phi_{m}$, its size, and the derived weight
$\sigma_{i}$ Inter-model variance of errors on sample ${\bm{x}}_{i}$
$\mathcal{D}^{*}$ Independent dataset for training the co-embedding extractor
${\bm{r}}_{m},{\bm{\mu}}_{c}\in\mathbb{R}^{D}$ Embedding of model $\phi_{m}$; embedding of channel $c$, ${\bm{\mu}}_{c}=\psi({\bm{x}}_{c})$
${\bm{R}}_{\text{zoo}}\in\mathbb{R}^{M\times D}$ Model zoo representation library of all TSFM embeddings
${\bm{R}}_{\text{task}}\in\mathbb{R}^{C\times D}$ Task representation matrix formed by embeddings of all $C$ channels